# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 4.2 DST model
The discrete self–trapping (DST) equation was introduced in [ELS] to describe the dynamics of small molecules, which then led to detailed studies of the DST dimer using different methods. The DST model we cast into our framework here is a generalisation of the dimer case to arbitrary (finite) degrees of freedom. This general case first appeared in [CJK] where its relationship with the periodic Toda chain was also hinted at. Our motivation here being different, we do not delve into this connection between the two theories. The interested reader is referred to [KSS] where this aspect was explored further.
Lax matrix:
We work here with the following avatar of the Lax matrix of the DST model
$L_{\text{DST}}(\lambda)=\dfrac{1}{\lambda}\sum_{i=1}^{T}c_{i}E_{ii}+\frac{1}{T}\sum_{i,j=1}^{T}\sum_{k=0}^{T-1}\frac{\omega^{k(j-i)}x_{i}X_{j}E_{ij}}{\lambda-\omega^{k}\zeta_{1}}+\sum_{i=1}^{T}E_{i,i+1}$ (4.25)
where $c_{i}$, for $i=1,\ldots,T$, are complex parameters, and $x_{i},X_{i}$ are the canonical coordinates satisfying the canonical Poisson bracket relations $\{X_{i},x_{j}\}=\delta_{ij}$, for $i,j=1,\ldots,T$, and the periodic conditions $(X_{0},x_{0})=(X_{T},x_{T})$ and $(X_{T+1},x_{T+1})=(X_{1},x_{1})$.
The DST Lax matrix in [KSS, Equation 3.8] is given as
$\widehat{L}_{\text{DST}}(\mu)=\sum_{i,j=1}^{T}\frac{b^{T+i-j}x_{i}X_{j}E_{ij}}{\mu-b^{T}}+\mu E_{T1}+\sum_{i\geq j}b^{i-j}x_{i}X_{j}E_{ij}+\sum_{i=1}^{T}c_{i}E_{ii}+\sum_{i=1}^{T-1}E_{i,i+1}$ (4.26)
where $b,c_{i}\in\mathbb{C}$ are parameters of the model, and $\mu$ is the spectral parameter of the Lax matrix. In the present setup where we realise the DST model as a cyclotomic Gaudin model, the role of the parameter $b$ is played by the location of the pole $\zeta_{1}$ on $\mathbb{C}P^{1}\setminus\{0,\infty\}$. One obtains the Lax matrix $L_{\text{DST}}(\lambda)$ by conjugating $\widehat{L}_{\text{DST}}(\lambda^{T})$ by the diagonal matrix $\mathcal{D}=\text{diag}(\lambda^{-1},\ldots,\lambda^{-T})$ followed by an overall multiplication by $\lambda^{-1}$, together with a change of $\lambda$-dependence, that is,
$L_{\text{DST}}(\lambda)=\lambda^{-1}\mathcal{D}\widehat{L}_{\text{DST}}(\lambda^{T})\mathcal{D}^{-1}\,,$ (4.27)
@@
The Poisson bracket of the Lax matrix $\widehat{L}_{\text{DST}}(\lambda)$ is given as
$\{\widehat{L}_{\text{DST}\,_{\mathbf{\underline{1}}}}(\lambda),\widehat{L}_{\text{DST}\,_{\mathbf{\underline{2}}}}(\mu)\}=[\widehat{r}_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),\widehat{L}_{\text{DST}\,_{\mathbf{\underline{1}}}}(\lambda)]-[\widehat{r}_{{}_{\mathbf{\underline{21}}}}(\mu,\lambda),\widehat{L}_{\text{DST}\,_{\mathbf{\underline{2}}}}(\mu)]$ (4.28)
@@
$\widehat{r}_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)=\frac{1}{\mu-\lambda}\left(\mu\sum_{i\leq j}+\lambda\sum_{i>j}\right)E_{ij}\otimes E_{ji}\,.$ (4.29)
Using the identity (4.1) once again, one finds that the gauge transformation (4.27) gives for $L_{\text{DST}}(\lambda)$ the $r$-matrix (4.8) associated with the cyclotomic $\mathfrak{gl}_{T}(\mathbb{C})$-Gaudin model, as we would have anticipated. We then have the Poisson bracket
$\{L_{\text{DST}\,_{\mathbf{\underline{1}}}}(\lambda),L_{\text{DST}\,_{\mathbf{\underline{2}}}}(\mu)\}=[r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),L_{\text{DST}\,_{\mathbf{\underline{1}}}}(\lambda)]-[r_{{}_{\mathbf{\underline{21}}}}(\mu,\lambda),L_{\text{DST}\,_{\mathbf{\underline{2}}}}(\mu)]\,.$ (4.30)
Similar to the case of the periodic Toda chain, the Lax matrix $L_{\text{DST}}(\lambda)$ can be seen as a realisation of the cyclotomic Gaudin Lax matrix, this time with simple poles at the origin and all $\omega^{k}\zeta_{1}$, $k\in\{0,\ldots,T-1\}$, for some $\zeta_{1}\in\mathbb{C}^{\times}$, and a double pole at infinity, that is,
$L_{\text{DST}}(\lambda)=\frac{K_{0}^{(0)}}{\lambda}+\frac{1}{T}\sum_{k=0}^{T-1}\frac{\sigma^{k}K_{1}}{\lambda-\omega^{k}\zeta_{1}}+K_{\infty}$ (4.31)
@@
$K_{0}^{(0)}=\sum_{i=1}^{T}c_{i}E_{ii}\,,\quad K_{1}=\sum_{i,j=1}^{T}x_{i}X_{j}E_{ij}\,,\quad K_{\infty}=\sum_{i=1}^{T}E_{i,i+1}\,.$ (4.32)
Orbit realisation:
Set $D=\{0,1,\infty\}$, and choose the non-dynamical element
$\Lambda_{\text{DST}}(\lambda)=\frac{\Lambda_{0}^{(0)}}{\lambda}+\frac{1}{T}\sum_{k=0}^{T-1}\frac{\sigma^{k}\Lambda_{1}}{\lambda-\omega^{k}\zeta_{1}}+\Lambda_{\infty}\in\Omega_{D}^{\Gamma}$ (4.33)
@@
$\Lambda_{0}^{(0)}=\sum_{i=1}^{T}c_{i}E_{ii}\in\mathfrak{g}^{(0)}\,,\quad\Lambda_{1}=E_{11}\in\mathfrak{g}\,,\quad\Lambda_{\infty}=\sum_{i=1}^{T}E_{i,i+1}\in\mathfrak{g}^{(1)}\,.$ (4.34)
The coadjoint action of the group elements $\varphi_{+}=\left(\varphi_{0+},\varphi_{1+},\varphi_{\infty+}\right)$, defined by (3.47), on $\Lambda_{\text{DST}}(\lambda)$ gives the orbit where $L_{\text{DST}}(\lambda)$ lives. From (3.55), we know that the components of $L_{\text{DST}}(\lambda)$ can then be expressed as
$K_{0}^{(0)}=\phi_{0}^{(0)}\Lambda_{0}^{(0)}\phi_{0}^{(0)\,-1}\,,\quad K_{1}=\phi_{1}\Lambda_{1}\phi_{1}^{-1}\,,\quad K_{\infty}=\Lambda_{\infty}\,,$ (4.35)
where we have denoted $\phi_{1}^{(0)}$ by $\phi_{1}$ for simplicity. This gives us a parameterisation of the Lax matrix $L_{\text{DST}}(\lambda)$ in (4.46) as an element of the coadjoint orbit ${\cal O}_{\Lambda}^{\text{DST}}$. We parameterise $\phi_{1}$ as
$\phi_{1}=\sum_{i,j=1}^{T}s_{ij}E_{ij}\,,$ (4.36)
and denote by $\hat{s}_{ij}$, $i,j=1,\ldots,T$, the entries of its inverse, that is,
$\phi_{1}^{-1}=\sum_{i,j=1}^{T}\hat{s}_{ij}E_{ij}\,.$ (4.37)
@@
$K_{0}^{(0)}=\sum_{i=1}^{T}c_{i}E_{ii}\,,\quad K_{1}=\sum_{i,j=1}^{T}s_{i1}\hat{s}_{1j}E_{ij}\,,\quad K_{\infty}=\sum_{i=1}^{T}E_{i,i+1}\,.$ (4.38)
@@
$x_{i}=s_{i1}\quad\text{and}\quad X_{i}=\hat{s}_{1i},\quad\text{for}\,\,i=1,\ldots,T\,,$ (4.39)
gives us the desired realisation of the Lax matrix $L_{\text{DST}}(\lambda)$ as an element of the coadjoint orbit ${\cal O}_{\Lambda}^{\text{DST}}$. Notice that $\operatorname{Tr}K_{1}=\sum_{i=1}^{T}x_{i}X_{i}=1$ is an orbit invariant and can be seen as being generated by the symmetry $x_{i}\rightarrow ax_{i}$, $X_{i}\rightarrow a^{-1}X_{i}$.
Lax equations:
Let us choose invariant functions $H_{p,r}$ on $\widetilde{\mathfrak{g}}_{D}^{(1)}$ as defined in (3.70), and treat $H_{p,0}$ and $H_{p,1}$ as the independent functions. The Lax equations for the DST model with respect to the elementary times $t_{p}^{r}$ are then given by
$\partial_{t_{p}^{r}}\iota_{\lambda}L_{\text{DST}}=[R_{\pm}\nabla H_{p,r}(\iota_{\lambda}L_{\text{DST}}),\iota_{\lambda}L_{\text{DST}}]\,.$ (4.40)
@@
$\partial_{t_{1}^{r}}L_{\text{DST}}=[M_{1,r},L_{\text{DST}}]\quad\text{with}\quad M_{1,r}=\begin{dcases}\quad 0&\text{for}\,\,r=0\\ -\dfrac{1}{T}\sum_{k=0}^{T-1}\dfrac{\omega^{k}\zeta_{1}\sigma^{k}K_{1}}{\lambda-\omega^{k}\zeta_{1}}&\text{for}\,\,r=1\,.\end{dcases}$ (4.41)
The $t_{1}^{0}$ equations are trivial. Similar to the case of the periodic Toda chain, the easiest way to get the $t_{1}^{1}$ equations of motion for $x_{i}$ and $X_{i}$, for $i=1,\ldots,T$, is to undo the dressing. With $M=K_{0}^{(0)}+\zeta_{1}K_{\infty}+\dfrac{1}{T}\sum_{k=1}^{T-1}\sigma^{k}K_{1}$, we find that $\phi_{1}^{-1}(\partial_{t_{1}^{1}}\phi_{1}\phi_{1}^{-1}-M)\phi_{1}$ must commute with $\Lambda_{1}$, and hence be block diagonal with a scalar “block”, say $\rho$, and a $(T-1)\times(T-1)$ block which is irrelevant. Then we find, collecting $x_{i}$ in the vector ${\bf x}$, and $X_{i}$ in the vector ${\bf X}$,
$\partial_{t_{1}^{1}}{\bf x}-M{\bf x}=\rho{\bf x},\quad\partial_{t_{1}^{1}}{\bf X}^{T}-{\bf X}^{T}M=\rho{\bf X}^{T}\,.$
Using the freedom $x_{i}\rightarrow ax_{i}$, $X_{i}\rightarrow a^{-1}X_{i}$ mentioned above with $a=e^{B}$, $\partial_{t_{1}^{1}}B=\rho$, we can set $\rho=0$. Explicitly, we obtain
$\begin{cases}\partial_{t_{1}^{0}}X_{i}=0\,,\\[6.45831pt] \partial_{t_{1}^{0}}x_{i}=0\,,\\[6.45831pt] \partial_{t_{1}^{1}}X_{i}=-c_{i}X_{i}-\zeta_{1}X_{i-1}-\dfrac{1}{T}\sum\limits_{k=1}^{T-1}\sum\limits_{j=1}^{T}\omega^{k(j-i)}x_{j}X_{j}X_{i}\,,\\[6.45831pt] \partial_{t_{1}^{1}}x_{i}=c_{i}x_{i}+\zeta_{1}x_{i+1}+\dfrac{1}{T}\sum\limits_{k=1}^{T-1}\sum\limits_{j=1}^{T}\omega^{k(j-i)}X_{j}x_{j}x_{i}\,.\end{cases}$ (4.42)
Lagrangian description:
We can now write a Lagrangian multiform for the DST hierarchy using Theorem 3.6 as
$\mathscr{L}_{\text{DST}}=\sum_{p=1}^{M}\mathscr{L}_{p,0}\mathrm{d}t_{p}^{0}+\sum_{p=1}^{M}\mathscr{L}_{p,1}\mathrm{d}t_{p}^{1}$ (4.43)
@@
$\mathscr{L}_{p,r}=\sum_{i=1}^{T}X_{i}\partial_{t_{p}^{r}}x_{i}-H_{p,r}(\iota_{\lambda}L_{\text{DST}}),\quad r\in\{0,1\}\,,$ (4.44)
where the potential term $H_{p,r}(\iota_{\lambda}L_{\text{DST}})$ is given by (3.70) for $r\in\{0,1\}$. Notice that we have dropped the kinetic contribution to $\mathscr{L}_{p,r}$ from the pole at origin since being a total derivative it will not enter the Euler-Lagrange equations. For $p=1$, the Lagrangian coefficients explicitly read
$\begin{split}&\mathscr{L}_{1,0}=\sum_{i=1}^{T}X_{i}\partial_{t_{1}^{0}}x_{i}-\dfrac{1}{2}\sum_{i=1}^{T}c_{i}^{2}\,,\\ &\mathscr{L}_{1,1}=\sum_{i=1}^{T}X_{i}\partial_{t_{1}^{1}}x_{i}-\dfrac{1}{2T}\sum_{i,j=1}^{T}\sum_{k=0}^{T-1}\omega^{k(j-i)}x_{i}x_{j}X_{i}X_{j}-\sum_{i=1}^{T}c_{i}x_{i}X_{i}-\zeta_{1}\sum_{i=1}^{T}x_{i+1}X_{i}\,.\end{split}$ (4.45)
It can be checked that the Euler-Lagrange equations obtained from varying $\mathscr{L}_{1,0}$ and $\mathscr{L}_{1,1}$ with respect to $X_{i},x_{i}$ are exactly the equations in (4.42).

### v2

@@ 4.2 DST model
The discrete self-trapping (DST) equation was introduced in [17] to describe the dynamics of small molecules, which then led to detailed studies of the DST dimer using different methods. The DST model we cast into our framework here is a generalisation of the dimer case to arbitrary (finite) degrees of freedom. This general case first appeared in [16] where its relationship with the periodic Toda chain was also hinted at. Our motivation here being different, we do not delve into this connection between the two theories. The interested reader is referred to [22] where this aspect was explored further.
Lax matrix. We work here with the following avatar of the Lax matrix of the DST model
$L_{{\rm DST}}(\lambda)=\dfrac{1}{\lambda}\sum_{i=1}^{T}c_{i}E_{ii}+\frac{1}{T}\sum_{i,j=1}^{T}\sum_{k=0}^{T-1}\frac{\omega^{k(j-i)}x_{i}X_{j}E_{ij}}{\lambda-\omega^{k}\zeta_{1}}+\sum_{i=1}^{T}E_{i,i+1},$
where $c_{i}$, for $i=1,\dots,T$, are complex parameters, and $x_{i}$, $X_{i}$ are the canonical coordinates satisfying the canonical Poisson bracket relations $\{X_{i},x_{j}\}=\delta_{ij}$ for $i,j=1,\dots,T$, and the periodic conditions $(X_{0},x_{0})=(X_{T},x_{T})$ and $(X_{T+1},x_{T+1})=(X_{1},x_{1})$.
The DST Lax matrix in [22, equation (3.8)] is given as
$\widehat{L}_{{\rm DST}}(\mu)=\sum_{i,j=1}^{T}\frac{b^{T+i-j}x_{i}X_{j}E_{ij}}{\mu-b^{T}}+\mu E_{T1}+\sum_{i\geq j}b^{i-j}x_{i}X_{j}E_{ij}+\sum_{i=1}^{T}c_{i}E_{ii}+\sum_{i=1}^{T-1}E_{i,i+1},$
where $b,c_{i}\in\mathbb{C}$ are parameters of the model, and $\mu$ is the spectral parameter of the Lax matrix. In the present setup where we realise the DST model as a cyclotomic Gaudin model, the role of the parameter $b$ is played by the location of the pole $\zeta_{1}$ on $\mathbb{C}P^{1}\setminus\{0,\infty\}$. One obtains the Lax matrix $L_{{\rm DST}}(\lambda)$ by conjugating $\widehat{L}_{{\rm DST}}\bigl{(}\lambda^{T}\bigr{)}$ by the diagonal matrix $\mathcal{D}={\rm diag}\bigl{(}\lambda^{-1},\dots,\lambda^{-T}\bigr{)}$ followed by an overall multiplication by $\lambda^{-1}$, together with a change of $\lambda$-dependence, that is,
$L_{{\rm DST}}(\lambda)=\lambda^{-1}\mathcal{D}\widehat{L}_{{\rm DST}}\bigl{(}\lambda^{T}\bigr{)}\mathcal{D}^{-1},$ (4.9)
@@
The Poisson bracket of the Lax matrix $\widehat{L}_{{\rm DST}}(\lambda)$ is given as
$\bigl{\{}\widehat{L}_{{\rm DST}\,_{\mathbf{\underline{1}}}}(\lambda),\widehat{L}_{{\rm DST}\,_{\mathbf{\underline{2}}}}(\mu)\bigr{\}}=\bigl{[}\widehat{r}_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),\widehat{L}_{{\rm DST}\,_{\mathbf{\underline{1}}}}(\lambda)\bigr{]}-\bigl{[}\widehat{r}_{{}_{\mathbf{\underline{21}}}}(\mu,\lambda),\widehat{L}_{{\rm DST}\,_{\mathbf{\underline{2}}}}(\mu)\bigr{]},$
@@
$\widehat{r}_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)=\frac{1}{\mu-\lambda}\biggl{(}\mu\sum_{i\leq j}+\lambda\sum_{i>j}\biggr{)}E_{ij}\otimes E_{ji}.$
Using the identity (4.1) once again, one finds that the gauge transformation (4.9) gives for $L_{{\rm DST}}(\lambda)$ the $r$-matrix (4.4) associated with the cyclotomic $\mathfrak{gl}_{T}(\mathbb{C})$-Gaudin model, as we would have anticipated. We then have the Poisson bracket
$\bigl{\{}L_{{\rm DST}\,_{\mathbf{\underline{1}}}}(\lambda),L_{{\rm DST}\,_{\mathbf{\underline{2}}}}(\mu)\bigr{\}}=\bigl{[}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),L_{{\rm DST}\,_{\mathbf{\underline{1}}}}(\lambda)\bigr{]}-\bigl{[}r_{{}_{\mathbf{\underline{21}}}}(\mu,\lambda),L_{{\rm DST}\,_{\mathbf{\underline{2}}}}(\mu)\bigr{]}.$
Similar to the case of the periodic Toda chain, the Lax matrix $L_{{\rm DST}}(\lambda)$ can be seen as a realisation of the cyclotomic Gaudin Lax matrix, this time with simple poles at the origin and all $\omega^{k}\zeta_{1}$, $k\in\{0,\dots,T-1\}$, for some $\zeta_{1}\in\mathbb{C}^{\times}$, and a double pole at infinity, that is,
$L_{{\rm DST}}(\lambda)=\frac{K_{0}^{(0)}}{\lambda}+\frac{1}{T}\sum_{k=0}^{T-1}\frac{\sigma^{k}K_{1}}{\lambda-\omega^{k}\zeta_{1}}+K_{\infty},$ (4.10)
@@
$K_{0}^{(0)}=\sum_{i=1}^{T}c_{i}E_{ii},\qquad K_{1}=\sum_{i,j=1}^{T}x_{i}X_{j}E_{ij},\qquad K_{\infty}=\sum_{i=1}^{T}E_{i,i+1}.$ (4.11)
Orbit realisation. Set $D=\{0,1,\infty\}$, and choose the non-dynamical element
$\Lambda_{{\rm DST}}(\lambda)=\frac{\Lambda_{0}^{(0)}}{\lambda}+\frac{1}{T}\sum_{k=0}^{T-1}\frac{\sigma^{k}\Lambda_{1}}{\lambda-\omega^{k}\zeta_{1}}+\Lambda_{\infty}\in\Omega_{D^{\prime}}^{\Gamma},$
@@
$\Lambda_{0}^{(0)}=\sum_{i=1}^{T}c_{i}E_{ii}\in\mathfrak{g}^{(0)},\qquad\Lambda_{1}=E_{11}\in\mathfrak{g},\qquad\Lambda_{\infty}=\sum_{i=1}^{T}E_{i,i+1}\in\mathfrak{g}^{(1)}.$
The coadjoint action of the group elements $\varphi_{+}=\left(\varphi_{0+},\varphi_{1+},\varphi_{\infty+}\right)$, defined by (3.24), on $\Lambda_{{\rm DST}}(\lambda)$ gives the orbit where $L_{{\rm DST}}(\lambda)$ lives. From (3.27), we know that the components of $L_{{\rm DST}}(\lambda)$ can then be expressed as
$K_{0}^{(0)}=\phi_{0}^{(0)}\Lambda_{0}^{(0)}\phi_{0}^{(0)\,-1},\qquad K_{1}=\phi_{1}\Lambda_{1}\phi_{1}^{-1},\qquad K_{\infty}=\Lambda_{\infty},$ (4.12)
where we have denoted $\phi_{1}^{(0)}$ by $\phi_{1}$ for simplicity. This gives us a parameterisation of the Lax matrix $L_{{\rm DST}}(\lambda)$ in (4.14) as an element of the coadjoint orbit ${\cal O}_{\Lambda}^{{\rm DST}}$. We parameterise $\phi_{1}$ as
$\phi_{1}=\sum_{i,j=1}^{T}s_{ij}E_{ij},$
and denote by $\hat{s}_{ij}$, $i,j=1,\dots,T$, the entries of its inverse, that is,
$\phi_{1}^{-1}=\sum_{i,j=1}^{T}\hat{s}_{ij}E_{ij}.$
@@
$K_{0}^{(0)}=\sum_{i=1}^{T}c_{i}E_{ii},\qquad K_{1}=\sum_{i,j=1}^{T}s_{i1}\hat{s}_{1j}E_{ij},\qquad K_{\infty}=\sum_{i=1}^{T}E_{i,i+1}.$
@@
$x_{i}=s_{i1}\qquad\text{and}\qquad X_{i}=\hat{s}_{1i}\qquad\text{for}\quad i=1,\dots,T,$
gives us the desired realisation of the Lax matrix $L_{{\rm DST}}(\lambda)$ as an element of the coadjoint orbit ${\cal O}_{\Lambda}^{{\rm DST}}$. Notice that $\operatorname{Tr}K_{1}=\sum_{i=1}^{T}x_{i}X_{i}=1$ is an orbit invariant and can be seen as being generated by the symmetry $x_{i}\rightarrow ax_{i}$, $X_{i}\rightarrow a^{-1}X_{i}$.
Lax equations. Let us choose invariant functions $H_{p,r}$ on $\widetilde{\mathfrak{g}}_{D}^{(1)}$ as defined in (3.36), and treat $H_{p,0}$ and $H_{p,1}$ as the independent functions. The Lax equations for the DST model with respect to the elementary times $t_{p}^{r}$ are then given by
$\partial_{t_{p}^{r}}\iota_{\lambda}L_{{\rm DST}}=[R_{\pm}\nabla H_{p,r}(\iota_{\lambda}L_{{\rm DST}}),\iota_{\lambda}L_{{\rm DST}}].$
@@
$\partial_{t_{1}^{r}}L_{{\rm DST}}=[M_{1,r},L_{{\rm DST}}]\qquad\text{with}\quad M_{1,r}=\begin{dcases}\quad 0&\text{for}\ r=0,\\ -\dfrac{1}{T}\sum_{k=0}^{T-1}\dfrac{\omega^{k}\zeta_{1}\sigma^{k}K_{1}}{\lambda-\omega^{k}\zeta_{1}}&\text{for}\ r=1.\end{dcases}$
The $t_{1}^{0}$ equations are trivial. Similar to the case of the periodic Toda chain, the easiest way to get the $t_{1}^{1}$ equations of motion for $x_{i}$ and $X_{i}$, for $i=1,\dots,T$, is to undo the dressing. With ${M=K_{0}^{(0)}+\zeta_{1}K_{\infty}+\frac{1}{T}\sum_{k=1}^{T-1}\sigma^{k}K_{1}}$, we find that $\phi_{1}^{-1}(\partial_{t_{1}^{1}}\phi_{1}\phi_{1}^{-1}-M)\phi_{1}$ must commute with $\Lambda_{1}$, and hence be block diagonal with a scalar “block”, say $\rho$, and a $(T-1)\times(T-1)$ block which is irrelevant. Then we find, collecting $x_{i}$ in the vector ${\bf x}$, and $X_{i}$ in the vector ${\bf X}$,
$\partial_{t_{1}^{1}}{\bf x}-M{\bf x}=\rho{\bf x},\qquad\partial_{t_{1}^{1}}{\bf X}^{T}-{\bf X}^{T}M=\rho{\bf X}^{T}.$
Using the freedom $x_{i}\rightarrow ax_{i}$, $X_{i}\rightarrow a^{-1}X_{i}$ mentioned above with $a={\rm e}^{B}$, $\partial_{t_{1}^{1}}B=\rho$, we can set ${\rho=0}$. Explicitly, we obtain
$\partial_{t_{1}^{0}}X_{i}=0,$
$\partial_{t_{1}^{0}}x_{i}=0,$
$\partial_{t_{1}^{1}}X_{i}=-c_{i}X_{i}-\zeta_{1}X_{i-1}-\dfrac{1}{T}\sum\limits_{k=1}^{T-1}\sum\limits_{j=1}^{T}\omega^{k(j-i)}x_{j}X_{j}X_{i},$
$\partial_{t_{1}^{1}}x_{i}=c_{i}x_{i}+\zeta_{1}x_{i+1}+\dfrac{1}{T}\sum\limits_{k=1}^{T-1}\sum\limits_{j=1}^{T}\omega^{k(j-i)}X_{j}x_{j}x_{i}.$ (4.13)
Lagrangian description. We can now write a Lagrangian multiform for the DST hierarchy using Theorem 3.6 as
$\mathscr{L}_{{\rm DST}}=\sum_{p=1}^{M}\mathscr{L}_{p,0}\,\mathrm{d}t_{p}^{0}+\sum_{p=1}^{M}\mathscr{L}_{p,1}\,\mathrm{d}t_{p}^{1}$
@@
$\mathscr{L}_{p,r}=\sum_{i=1}^{T}X_{i}\partial_{t_{p}^{r}}x_{i}-H_{p,r}(\iota_{\lambda}L_{{\rm DST}}),\qquad r\in\{0,1\},$
where the potential term $H_{p,r}(\iota_{\lambda}L_{{\rm DST}})$ is given by (3.36) for $r\in\{0,1\}$. Notice that we have dropped the kinetic contribution to $\mathscr{L}_{p,r}$ from the pole at origin since being a total derivative it will not enter the Euler–Lagrange equations. For $p=1$, the Lagrangian coefficients explicitly read
$\mathscr{L}_{1,0}=\sum_{i=1}^{T}X_{i}\partial_{t_{1}^{0}}x_{i}-\dfrac{1}{2}\sum_{i=1}^{T}c_{i}^{2},$
$\mathscr{L}_{1,1}=\sum_{i=1}^{T}X_{i}\partial_{t_{1}^{1}}x_{i}-\dfrac{1}{2T}\sum_{i,j=1}^{T}\sum_{k=0}^{T-1}\omega^{k(j-i)}x_{i}x_{j}X_{i}X_{j}-\sum_{i=1}^{T}c_{i}x_{i}X_{i}-\zeta_{1}\sum_{i=1}^{T}x_{i+1}X_{i}.$
It can be checked that the Euler–Lagrange equations obtained from varying $\mathscr{L}_{1,0}$ and $\mathscr{L}_{1,1}$ with respect to $X_{i}$, $x_{i}$ are exactly the equations in (4.13).